OpenStudy (anonymous):

Find the area of the sector formed by central angle (theta) = 3pi/5 in a circle of radius r=4m.

OpenStudy (lalaly):

\[\underline{\text{area of sector}}=\frac{1}{2}r^2 \theta\]\[=\frac{1}{2}(4m)(\frac{3}{5}\pi)\]

OpenStudy (anonymous):

So, would I be left with 2m (3/5pi)? Then, left with 6pi/5?

OpenStudy (lalaly):

the m is meters? or just a variable?

OpenStudy (anonymous):

It's meters.

OpenStudy (lalaly):

oh then yeah u got it:)

OpenStudy (anonymous):

Well, when I put it in the calculator, it came out to be 3.77. That answer is wrong somehow. :/

OpenStudy (lalaly):

why wrong? ..do u have the right answer?

OpenStudy (anonymous):

I'm pretty sure I did it right, by the formula you gave me. That is a correct formula.

OpenStudy (lalaly):

ohhhh i did something wrong lol

OpenStudy (lalaly):

forgot to square the radius\[A=\frac{1}{2}(4)^2(\frac{3}{5}\pi)\]

OpenStudy (lalaly):

Im sorry -_-'

OpenStudy (anonymous):

Oh!!! No, it's fine!

OpenStudy (lalaly):

Now you'll get it right:)

OpenStudy (anonymous):

Mhm!!! :D

OpenStudy (anonymous):

Got it! :) Thank you!!

OpenStudy (lalaly):

:)